I am suppose to write an equation for the following, but I am not sure I am doing it right
Rboert buys 6 tomato plants at a certain price, he needs $1.00 more than he needs to buy 7 plants at $0.20 less per plant.

Answer by ptaylor(2198)


Lets start by dealing in pennies, that will make life a little easier. (1) It gets rid of the decimals and (2) It reduces the chance of getting pennies and dollars confused.
Let x=price he paid for each of the 6 plants
Then 6x=cost of all six plants
And x-20 is the cost of each of the 7 plants
Now we are told that he needs 100 (pennies) more (when he buys the 6 plants) (that would be 6x+100) than he needs to buy 7 plants at 20 cents less per plant (that would be 7(x-20). so our equation to solve is:
6x+100=7(x-20) get rid of parens--------------this is the equation
6x+100=7x-140 subtract 7x and also 100 from both sides
6x-7x+100-100=7x-7x-140-100 collect like terms
-x=-240 multiply both sides by -1
x=240 pennies or
x=$2.40 per plant for the 6 plants
x-20=240-20=220 or
x-20=$2.20 cost per plant of the 7 plants
CK
6*240+100=7*220
1440+100=1540
1540=1540
Hope this helps--------------------ptaylor